OpenStudy (anonymous):

How many solutions does the system of equations below have? 2y - x = 5 3x - 6y = 5

OpenStudy (anonymous):

??? what would the answer be?

OpenStudy (anonymous):

do you just need to solve for y?

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

well i go y=1/2x+5/2

OpenStudy (anonymous):

how many solutions does it have??

OpenStudy (anonymous):

id say 2

OpenStudy (anonymous):

so it would be infinite number??correct?

OpenStudy (anonymous):

ya that sounds right.

OpenStudy (anonymous):

How many solutions does this system of equations have? x+y= -4 x -y= 0

OpenStudy (anonymous):

what about that problem how many solutions??

OpenStudy (anonymous):

still four total solutions if your solving for both x and y

OpenStudy (anonymous):

so it would be infinite number correct?

OpenStudy (anonymous):

im not exactly sure what you mean by that